# Small Language Models can Outperform Humans in Short Creative Writing: A Study Comparing SLMs with Humans and LLMs

### v1

@@ Appendix B Assessors Information
We recruited 68 volunteer participants for our study. All subjects recruited for our experimentation were students of an international Master in Management. All students were proficient in English (which is the official language of the master), although they come from different countries and therefore have different cultural backgrounds. They also have different academic backgrounds (sciences and technology, social sciences and humanities). Our set of assessors is, therefore, quite homogeneous in terms of age and educational level, but quite diverse in cultural and educational background.
@@ C.2 Linear Mixed-Effects Models
Linear Mixed-Effects Models (Bates et al., 2014; Kaptein, 2016) is a statistical method that consist of a linear regression where the model also includes random effects to account for the variance produced by the subjects and the set of the items (in our case, synopses) selected for the experiment. We have performed a Linear Mixed-Effects Models regression with two goals in mind. First, we want to verify the results we observed in the previous sections with a more in-depth statistical analysis. And, secondly, we want to check if the subject profiles influence their assessments.
@@
In conclusion, the results of the first linear mixed effect model show that subject attributes (age, language and gender) do not have a significant impact in the responses. The second model confirms the significance results obtained with the Wilcoxon signed-rank test in the body of the article.
@@ C.3.1 Reader Outliers
Using these criteria, seven readers where marked as outliers. Eliminating them does not change the results significantly, and in fact it increases the difference in favor of the SLM as Table 8 shows. Therefore, given the difficulty of differentiating between a true outlier and a random outlier, we decided to keep all participants for our analysis.
@@ C.3.2 Synopses Outliers
During the qualitative evaluation shown below, we detected 2 human synopses that seemingly contained subtle human errors. Removing them makes the average quality of human synopsis grow. However, the mean of SLM-generated synopses remains higher even removing these two faulty human synopses.
Appendix D Linguistic Analysis BAT vs. GPT 3.5 and GPT-4o

### v2

@@ Appendix B Assessors Information
We recruited 68 volunteer participants for our study. All subjects recruited for our experimentation were students of an international Master’s program in Management. All students were proficient in English (the official language of the Master’s program), although they come from different countries and therefore have different cultural backgrounds. They also have different academic backgrounds (sciences and technology, social sciences and humanities). Our set of assessors is, therefore, relatively homogeneous in terms of age and current educational level (all are Master’s students), but quite diverse in cultural and academic backgrounds.
@@ C.2 Linear Mixed-Effects Models
Linear Mixed-Effects Models (Bates et al., 2014; Kaptein, 2016) are a statistical method that consist of a linear regression where the model also includes random effects to account for the variance produced by the subjects and the set of the items (in our case, synopses) selected for the experiment. We have performed a Linear Mixed-Effects Models regression with two goals in mind. First, we want to verify the results we observed in the previous sections with a more in-depth statistical analysis. And, secondly, we want to check if the subject profiles influence their assessments.
@@
In conclusion, the results of the first linear mixed effect model show that subject attributes (age, educational background, and language) do not have a significant impact in the responses. The second model confirms the significance results obtained with the Wilcoxon signed-rank test in the body of the article.
@@ C.3.1 Reader Outliers
Using these criteria, seven readers where marked as outliers. Eliminating them does not change the results significantly. In fact, it increases the difference in favor of the SLM as Table 8 shows. Therefore, given the difficulty of differentiating between a true outlier and a random outlier, we decided to keep all participants for our analysis.
@@ C.3.2 Synopses Outliers
During the qualitative evaluation shown below, we detected two human synopses that seemingly contained subtle human errors. Removing them makes the average quality of human synopsis grow. However, the mean of SLM-generated synopses remains higher even removing these two faulty human synopses.
Appendix D Linguistic Analysis BART-large vs. GPT 3.5 and GPT-4o